OpenStudy (anonymous):

Hello everyone, i am having trouble with school work.again..not a surprise.... Any way here is the word problem: Rocky McFist spends a total of 90$/week for karate and ballet lessons.If the amount spend on karate was increased by 10 $/week, it will be 2/3s the of the amount spent on ballet lessons.How much does rocky spend on ballet lessons.

OpenStudy (anonymous):

can anybody help?

OpenStudy (anonymous):

Let x = price of ballet and y = price of karate x+y=90 y+10 = 2/3 x

OpenStudy (anonymous):

solve the second one for karate y = 2/3x - 10 substitute into the first equation x + 2/3x - 10 = 90 5/3 x = 100 x = 60 Ballet lessons are $60 a week

OpenStudy (anonymous):

oh kay

OpenStudy (anonymous):

could you help me with another one?

OpenStudy (anonymous):

I can, post it to the right so if I get busy with other things others can help.

OpenStudy (anonymous):

okay